OpenStudy (anonymous):

How to convert 1 OVER square root of 5 to an exponential expression??

OpenStudy (anonymous):

GO GO GO :] Ty in advance XD

OpenStudy (anonymous):

\[\sqrt{b}=b^{\frac{1}{2}}\] and so \[\frac{1}{\sqrt{5}}=5^{-\frac{1}{2}}\]

OpenStudy (anonymous):

Can explain plz XD